The width of a rectangle is 9 cm shorter than its length. The peremeter of the rectangle is 40 cm. What are the longth and with of the rectangle?
Answer by Alan3354(69443)

Do it like this one:
----------------------
The length of a rectangle is 10 cm longer than its width. The perimeter of the rectangle is 68 cm. What are the length and width of the rectangle?
-------
Perimeter = 2W + 2L = 68
W + L = 34
L = W + 10
Sub for L in the 1st eqn
W + (W+10) = 34
2W + 10 = 34
W = 12 cm
L = 22 cm